OpenStudy (anonymous):

determine the sum of the first six terms 2-6+18;....

Directrix (directrix):

@Michy --> What's the pattern for the terms? I see that the signs seem to alternate. With three terms, it's dicey, for me at least.

OpenStudy (anonymous):

That question came directly from the exercise, the way I did it is 2-6+18= 14, 2-6+14=10, 2-6+10=6, 2-6+6=2, 2-6+2=-2, 2-6-2=-6 =14+10+6+2+(-2)+(-6) the constant is 4 =24 I don't know... I'm usually wrong O_o

Directrix (directrix):

I think the pattern is to multiply the previous term by -3 (negative 3) after being given the first term. So I got the following: 3 - 6 + 18 -54 +172 - 516. Will you check those and add them? Post your result here.